Pleas help me solve and understand the this problem: A biologist is studying a fish. she uses the fishes length in cm L to estimate the fishes age in years T according to the formula T=19-5ln(53-L)
a) estimate the age of a 35 cm fish
b) according to the formula what length should a 10yr old fish be.
pleas include steps when solving the problem.

Answer by rothauserc(4718)

T = 19 - 5ln(53-L)
:
a) T = 19 - 5ln(53-35)
:
T = 19 - 5ln(18)
:
T = 19 - 5 * 2.8904 = 4.548
:
*****************************************
35 cm fish is approximately 4.5 years old
*****************************************
:
b) 10 = 19 - 5ln(53-L)
:
5ln(53-L) = 19 - 10 = 9
:
ln(53-L) = 9/5 = 1.8
:
53-L = inverse log(base e) of 1.8 = 6.0496
:
L = 53 - 6.0496 = 46.9504 approximately 47
:
*****************************************
A 10 year old fish should be 47 cm long
*****************************************
: